Perdon Corporation manufactures safes—large mobile safes, and large walk-in stationary bank safes. As part of its annual budgeting process, Perdon is analyzing the profitability of its two products. Part of this analysis involves estimating the amount of overhead to be allocated to each product line. The information shown below relates to overhead.

Mobile Safes

Walk-in Safes

Units planned for production 210 51
Material moves per product line 310 250
Purchase orders per product line 450 340
Direct labor hours per product line 810 1,710

The total estimated manufacturing overhead was $276,000. Under traditional costing (which assigns overhead on the basis of direct labor hours), what amount of manufacturing overhead costs are assigned to: (Round answers to 2 decimal places, e.g. 12.25.)

(1) One mobile safe

$enter a dollar amount per unit rounded to 2 decimal places

per unit
(2) One walk-in safe

$enter a dollar amount per unit rounded to 2 decimal places

per unit

eTextbook and Media

  

  

The total estimated manufacturing overhead of $276,000 was comprised of $163,000 for materials handling costs and $113,000 for purchasing activity costs. Under activity-based costing (ABC): (Round answers to 2 decimal places, e.g. 12.25.)

What amount of materials handling costs are assigned to:

(a) One mobile safe

$enter a dollar amount per unit rounded to 2 decimal places

per unit
(b) One walk-in safe

$enter a dollar amount per unit rounded to 2 decimal places

per unit

eTextbook and Media

  

  

The total estimated manufacturing overhead of $276,000 was comprised of $163,000 for materials handling costs and $113,000 for purchasing activity costs. Under activity-based costing (ABC): (Round answers to 2 decimal places, e.g. 12.25.)

What amount of purchasing activity costs are assigned to:

(a) One mobile safe

$enter a dollar amount per unit rounded to 2 decimal places

per unit
(b) One walk-in safe

$enter a dollar amount per unit rounded to 2 decimal places

per unit

eTextbook and Media

  

  

Compare the amount of overhead allocated to one mobile safe and to one walk-in safe under the traditional costing approach versus under ABC. (Round answers to 2 decimal places, e.g. 12.25.)

Traditional Costing

Activity-Based Costing

Mobile safe

$enter a dollar amount rounded to 2 decimal places

$enter a dollar amount rounded to 2 decimal places

Walk-in safe

$enter a dollar amount rounded to 2 decimal places

$enter a dollar amount rounded to 2 decimal places

eTextbook and Media

Saddle Inc. has two types of handbags: standard and custom. The controller has decided to use a plantwide overhead rate based on direct labor costs. The president has heard of activity-based costing and wants to see how the results would differ if this system were used. Two activity cost pools were developed: machining and machine setup. Presented below is information related to the company’s operations.

Standard

Custom

Direct labor costs $51,000 $116,000
Machine hours 1,410 1,330
Setup hours 96 400


Total estimated overhead costs are $304,000. Overhead cost allocated to the machining activity cost pool is $195,000, and $109,000 is allocated to the machine setup activity cost pool.

Incorrect answer iconYour answer is incorrect.

Compute the overhead rate using the traditional (plantwide) approach. (Round answer to 2 decimal places, e.g. 12.25.)

Predetermined overhead rate

enter the predetermined overhead rate in percentages of direct labor cost rounded to 2 decimal places

% of direct labor cost

eTextbook and Media

  

  

Correct answer iconYour answer is correct.

Compute the overhead rates using the activity-based costing approach. (Round answers to 2 decimal places, e.g. 12.25.)

Machining

$enter a dollar amount per machine hour rounded to 2 decimal places

per machine hour
Machine setup

$enter a dollar amount per setup hour rounded to 2 decimal places

per setup hour

eTextbook and Media

  

  

Incorrect answer iconYour answer is incorrect.

Determine the difference in allocation between the two approaches. (Round answers to 0 decimal places, e.g. 1,225.)

Traditional costing
Standard

$enter a dollar amount rounded to 0 decimal places

Custom

$enter a dollar amount rounded to 0 decimal places

Activity-based costing
Standard

$enter a dollar amount rounded to 0 decimal places

Custom

$enter a dollar amount rounded to 0 decimal places

eTextbook and Media
Answer #1

Overhead per hour Estimated overheads/Estimated hours

= 276000/(810+1710)

= $109.52 per hour

One Mobile safe = 109.52*810/210= $422.43 per unit

One walk in safe = 109.52*1710/51 = $3672.14 per unit

Handling cost

Mobile Safe = (163000*310/560)/210 = $429.68 per unit

Walk in = (163,000*250/560)/51 = $1426.82 per unit

Purchase activity

Mobile = (113000*450/790)/210 = 306.51 per unit

Walk in = (113000*340/790)/51 = $953.59 per unit

Traditional

ABC

Mobile Safe

422.43

736.19

Walk-in Safe

3672.14

2380.41